OpenStudy (anonymous):

Please help . Only 8th grade ques

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

Plz wait for 5 mins

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

i might be able to help too

OpenStudy (anonymous):

s = ut + 1 at2 (2 is square) ------ 2

OpenStudy (anonymous):

Then what is the value of u

OpenStudy (anonymous):

Plz help

OpenStudy (michele_laino):

are you asking for inverse formula?

OpenStudy (texaschic101):

is the problem : s = ut + 1/2^2 ??

OpenStudy (anonymous):

No it is s = 1/2 at^2

OpenStudy (anonymous):

And u have to change the subject to u. So what is the value for 'u'